OpenStudy (anonymous):

Write the equation of the line given the slope and the y-intercept. Type your answer in the box without skipping any spaces. m = 2, y-intercept (0, 3)

OpenStudy (katrinakaif):

Put that into the equation: (y -y1) = m (x-x1) as (0,3) = (x1,y1) and m = 2 You should get your equation =)

OpenStudy (anonymous):

(y-3)= 2(x-0)

OpenStudy (anonymous):

is that right?

OpenStudy (katrinakaif):

Yes =) Now distribute the 2 and simplify